# Wi-CBR: WiFi-based Cross-domain Behavior Recognition via Multimodal Collaborative Awareness

### v1

@@ Abstract
WiFi-based human behavior recognition aims to recognize gestures and activities by analyzing wireless signal variations. However, existing methods typically focus on a single type of data, neglecting the interaction and fusion of multiple features. To this end, we propose a novel multimodal collaborative awareness method. By leveraging phase data reflecting changes in dynamic path length and Doppler Shift (DFS) data corresponding to frequency changes related to the speed of gesture movement, we enable efficient interaction and fusion of these features to improve recognition accuracy. Specifically, we first introduce a dual-branch self-attention module to capture spatial-temporal cues within each modality. Then, a group attention mechanism is applied to the concatenated phase and DFS features to mine key group features critical for behavior recognition. Through a gating mechanism, the combined features are further divided into PD-strengthen and PD-weaken branches, optimizing information entropy and promoting cross-modal collaborative awareness. Extensive in-domain and cross-domain experiments on two large publicly available datasets, Widar3.0 and XRF55, demonstrate the superior performance of our method.
1 Introduction
Human Behavior Recognition (HBR) plays a crucial role in applications such as smart surveillance, healthcare, smart homes, and human-computer interaction. HBR can be classified into sensor-based, vision-based, and wireless signal-based [1, 2, 3] approaches. With the widespread deployment of WiFi systems and advancements in wireless technology, HBR using wireless sensing has gained significant attention. HBR scenarios span indoor environments like homes, workplaces, and medical facilities, while behavior types include human-object, human-human, and human-computer interaction. Coarse-grained activities can be easily detected using motion tracking techniques, but fine-grained gestures, which involve subtle movements and minimal signal disruption, remain challenging to recognize [4].
WiFi signals, characterized by Channel State Information (CSI) or Received Signal Strength Indicator (RSSI), are used to model the wireless channel. The core idea of WiFi sensing is that the changes in CSI patterns due to human or object motion contain unique information. Gesture recognition, for example, is achieved by learning the mapping between these signal variations and human behaviors [5, 6]. CSI-based gesture recognition for cross-domain tasks can be broadly categorized into modeling-based and learning-based approaches.
Modeling-Based HBR. Model-based approaches primarily focus on developing signal processing algorithms to extract environment-independent features from received signals, such as Body Coordinate Velocity Profile (BVP) [7], Motion Change Pattern (MCP) [8], and Doppler Frequency Shift (DFS) [9]. However, these methods have limitations, particularly in complex gesture recognition tasks, due to their inability to capture the full range of original data, which leads to reduced performance.
Learning-Based HBR. Recent advancements in deep learning have revolutionized various fields [10, 11], including Wi-Fi sensing. However, CSI signals present a significant challenge for learning-based approaches due to their sensitivity to environmental factors, such as the user’s location, orientation, and the surrounding environment. The received CSI signals not only encode information about the target object but also capture environment-specific details, such as static paths and noise, that are unrelated to the object being sensed [12, 13]. In cross-domain recognition, where the distribution of training and test data differs, this environmental noise complicates the task, necessitating the use of domain-invariant features to guide the learning process.
Figure 1: Wi-CBR Processing Flow.
@@ 1 Introduction
Due to the inherent limitations of the above approaches, we propose Wi-CBR to leverage the full spectrum of raw data while ensuring the integration of domain-independent features Fig 1. Specifically, our method introduces a novel multimodal collaborative awareness framework that efficiently combines phase data, which captures dynamic path length changes, and Doppler Shift (DFS) data, which reflects frequency shifts tied to gesture speed. To capture spatial-temporal patterns within each modality, we employ a dual-branch self-attention module, enabling the system to focus on important temporal and spatial features within each signal type. A group attention mechanism is then applied to the concatenated phase and DFS features, allowing the model to identify key group features that are essential for behavior recognition. Finally, a gating mechanism is used to divide the fused features into enhancement and suppression branches, optimizing information entropy and facilitating cross-modal collaborative complementarity. This fusion of multiple data sources, along with the innovative use of attention and gating mechanisms, enables more accurate and robust behavior recognition, particularly in cross-domain scenarios. Our main contributions are summarized as follows:
  • We propose a multimodal collaborative awareness method for WiFi-based behavior recognition, which exploits two modal cues: phase and DFS. Phase reflects dynamic path length changes, while DFS captures the frequency variations related to gesture movement speed.
  • A two-branch self-attention learning module is developed to capture the spatiotemporal cues within the modality, and a group attention mechanism is designed to mine key group features from the concatenated phase and DFS data.
  • Through a gating mechanism, combined features are divided into PD-strengthen and PD-weaken branches, and collaborative awareness interactions are further performed to improve cross-modal complementarity.
2 Related Work
WiFi-based human behavior recognition [14, 15] can be broadly classified into modeling-based and learning-based approaches. Model-based methods rely on specialized knowledge to extract context-independent features from the original signal, but this can lead to the loss of important information. On the other hand, learning-based methods use machine learning for gesture recognition, though they often lack domain-independent feature guidance, limiting their effectiveness.
2.1 Modeling-Based HBR
Modeling-based approaches involve preprocessing raw CSI data, extracting manual features (e.g., velocity distribution statistics), and using machine learning models for gesture recognition [6, 16, 17, 18, 19, 20, 21, 22].
WiFinger [6] uses detailed CSI to recognize sign language, mitigating environmental noise and adapting to gesture inconsistency. Gao et al. [18] modeled the relationship between gesture signals and noise, using a dynamic phase exponential error (edp exponent) for gesture quality. Wi-NN [22] applies time-domain feature selection and a KNN classifier for recognition. WiGest [5] uses manually constructed patterns, but its reliance on coarse-grained RSS limits accuracy. WiMU [23] faces scalability issues due to manual pattern definitions, while WiDraw [24] requires over 25 WiFi transceivers for accurate hand tracking, making it impractical. QGesture [25] recognizes gestures with two antennas, but requires prior knowledge of hand position.
While these studies establish a correlation between WiFi signals and gestures, they overlook the impact of different environments on gesture recognition. Performing the same gesture in different environments can significantly alter WiFi waveforms [26, 27]. WiDar3 [7] introduces a domain-agnostic feature, BVP, unaffected by environmental factors, while WiHF [8] derives domain-independent motion patterns for cross-domain recognition. However, handcrafted features limit full spatiotemporal cue capture. WiGesture [28] shifts focus to hand-oriented features, extracting Motion Navigation Primitives (MNP) independent of position.
2.2 Learning-Based HBR
Learning-based methods rely on directly learning raw CSI data, such as amplitude and phase, for automatic pattern recognition.
Wikey [29] applies WiFi sensing for keystroke recognition, but it is highly sensitive to environmental changes. WiSign [30] extracts spatio-temporal features of amplitude and phase for continuous sign language recognition, but deep learning-based approaches require extensive domain-specific training. Tong et al. [31] developed a CSI-based gesture recognition system using a dynamic CNN-GRU-Attention (CGA) model and introduced techniques like phase correction and adaptive gesture action truncation to enhance data validity. Yang et al. [32] proposed a deep concatenated representation learning architecture combining CNNs and bidirectional RNNs for one-time gesture recognition, improving spatiotemporal pattern learning. WiHGR [33] builds a phase difference matrix and uses an improved Attention-Based Bidirectional Gate Recurrent Unit (ABGRU) to automatically extract discriminative features.
To address cross-domain challenges, some learning-based methods have also emerged. ABLSTM [34] introduced an attention-based framework that avoids the loss of implicit knowledge caused by hand-crafted features. CROSSGR [35] extracts user-independent, gesture-related features from WiFi channel data, enhancing the system’s utility without prior knowledge of the user. WiGr [36] utilizes query sample and class prototype similarity for gesture classification, mitigating cross-domain CSI variations. WIGRUNT [37] proposed a spatio-temporal dual-attention network that processes phase data into images for feature extraction using ResNet. Wi-SFDAGR  [38]is the most recent work that tackles the cross-domain problem using Unsupervised Domain Adaptation (UDA), addressing scenarios where source data is unavailable during adaptation to unlabeled test data.
3 Method
Figure 2: Phase and DFS visualization.
@@ 3 Method
Wi-CBR is divided into four main parts: signal preprocessing, two-branch intra-modal self-attention learning, cross-modal collaborative interaction fusion, and classification prediction. In signal preprocessing, the CSI-ratio model is used for denoising,and the DFS is obtained using STFT. Subsequently, the phase and DFS matrices of the CSI ratios are visualized as images suitable for deep learning processing. In the network part, initial feature extraction is achieved by two-branch intra-modal self-attentive learning as well as pre-trained ResNet18. Then, the feature fusion is achieved by a cross-model interactive fusion module. Finally, the behavior prediction is achieved by a classifier, and the two kinds of losses achieve constraints.As shown in Fig 3.
3.1 Task Definition
@@
$H(f,t)=H_{s}(f,t)+H_{d}(f,t)=\sum_{m_{s}\in\Phi_{s}}a_{m_{s}}(f,t)e^{-j2\pi\frac{d_{m_{s}}(t)}{\lambda}}+\sum_{m_{d}\in\Phi_{d}}a_{m_{d}}(f,t)e^{-j2\pi\frac{d_{m_{d}}(t)}{\lambda}}.$ (3)
3.2 CSI Denoising Preprocessing
@@
As demonstrated in the previous section, the gesture can be portrayed by the change of phase-shift in CSI. Unfortunately, for commodity WiFi devices, as the transmitter and receiver are not synchronized, there exists a time-varying random phase offset $e^{-j\theta_{\text{offset}}}$:
$H(f,t)=e^{-j\theta_{\text{offset}}}(H_{s}(f,t)+H_{d}(f,t))=e^{-j\theta_{\text{offset}}}\left(H_{s}(f,t)+A(f,t)e^{-j2\pi\frac{d(t)}{\lambda}}\right),$ (4)
@@
where $A(f,t)$, $e^{-j2\pi\frac{d(t)}{\lambda}}$, and $d(t)$ denote the complex attenuation, phase-shift, and path length of dynamic components, respectively. This random offset thus prevents us from directly using the CSI phase information.
@@
Therefore, we need to eliminate $e^{-j\theta_{\text{offset}}}$. Fortunately, for commodity WiFi cards, this random offset remains the same across different antennas on the same WiFi network interface card (NIC) as they share the same RF oscillator. Thus, it can be eliminated by the CSI-ratio model [16, 25]:
$H_{r}(f,t)=\frac{H_{1}(f,t)}{H_{2}(f,t)}=\frac{e^{-j\theta_{\text{offset}}}(H_{s,1}+H_{d,1})}{e^{-j\theta_{\text{offset}}}(H_{s,2}+H_{d,2})}=\frac{A_{1}e^{-j2\pi\frac{d_{1}(t)}{\lambda}}+H_{s,1}}{A_{2}e^{-j2\pi\frac{d_{1}(t)+\Delta d}{\lambda}}+H_{s,2}},$ (5)
@@
where $H_{1}(f,t)$ and $H_{2}(f,t)$ are the CSI of two receiving antennas. When two antennas are close to each other, $\Delta d$ can be regarded as a constant. According to Mobius transformation  [39], (5) represents transformations such as scaling and rotation of the phase-shift $e^{-j2\pi\frac{d_{1}(t)}{\lambda}}$ of antenna 1 in the complex plane, and these transformations will not affect the changing trend of the phase-shift[40, 41, 42].
CSI to Phase. Phase $P$ extracted from $H_{r}$ can be used to describe gestures:
$Phase=angle(H_{r}),$ (6)
@@
where $angle(\cdot)$ denotes the phase extraction function. For a complex $z=\text{abs}(z)\cdot e^{j\theta}$, we can use $angle(\cdot)$ to obtain the phase of $z$, $\theta=angle(z)$.The visualized Phase image is shown in Fig. 2(a).
CSI to DFS. The DFS captures frequency changes due to gesture motion, such as hand speed. To extract DFS, the Short-Time Fourier Transform is applied to $H_{q}(f,t)$:
[…]
@@
where: $w(t-\tau)$: Window function (e.g., Hanning window) to segment the signal. $\tau$: Time localization. $\omega$: Frequency in radians, tied to the Doppler shift. The power spectrogram is then computed:
[…]
@@
In this spectrogram:The frequency $\omega$ at each time $\tau$ corresponds to the Doppler shift $f_{d}=\frac{\omega}{2\pi}$.$f_{d}$ is proportional to the gesture speed $v(t)$.Positive frequencies indicate motion toward the receiver, while negative frequencies indicate motion away.This time-frequency representation enables detailed analysis of gesture dynamics.For the Widar3.0 dataset, each DFS file is a 6*121*T matrix, where the first dimension represents the 6 receivers, the second dimension represents the 121 frequency segments ranging from [-60, 60] Hz, and the third dimension represents the timestamps with a sampling rate of 1000Hz. The visualized DFS image is shown in Fig. 2(b).
Figure 3: The System framework of Wi-CBR. Wi-CBR is divided into four main parts: Signal Preprocessing, Two-branch Intra-modal Self-attention Learning, Cross-modal Collaborative Interaction Fusion, and Gesture Recognition.
3.3 Two-branch Intra-modal Self-attention Learning
@@
$\mathbf{A}_{p}$ $=\sigma\left(\text{BN}\left(f_{\text{conv}}^{(7)}\left(f_{\text{conv}}^{(7)}(\mathbf{P}_{x})\right)\right)\right);$ $\quad\mathbf{A}_{d}$ $=\sigma\left(\text{BN}\left(f_{\text{conv}}^{(7)}\left(f_{\text{conv}}^{(7)}(\mathbf{D}_{x})\right)\right)\right),$ (9)
where $\sigma(\cdot)$ is the sigmoid function. The refined features are obtained through:
$\mathbf{P}_{\text{out}}$ $=\mathbf{P}_{x}\otimes\mathbf{A}_{p}\oplus\mathbf{P}_{x};$ $\quad\mathbf{D}_{\text{out}}$ $=\mathbf{D}_{x}\otimes\mathbf{A}_{d}\oplus\mathbf{D}_{x},$ (10)
where $\otimes$ is element-wise multiplication, $\oplus$ is element-wise summation. Each branch maintains independent convolution parameters, with channel dimensions preserved in first convolution ($3\rightarrow 3$) and reduced to single channel ($3\rightarrow 1$) in the second convolution.
@@
$\mathbf{X}_{P}&=\phi_{\text{ResNet}}^{P}(\mathbf{P}_{\text{out}});&\quad\mathbf{X}_{D}&=\phi_{\text{ResNet}}^{D}(\mathbf{D}_{\text{out}})\end{aligned},$ (11)
where $\phi_{\text{ResNet}}^{P}$ and $\phi_{\text{ResNet}}^{D}$ denote independent ResNet-18 backbones. The final representation combines features through:
$\mathbf{X}_{\text{PD}}=\text{Concat}(\mathbf{X}_{P},\mathbf{X}_{D})\in\mathbb{R}^{1024\times 7\times 7}.$ (12)
3.4 Cross-modal Synergistic Interaction and Fusion
@@
Cross-modal Interaction. To make the extracted Phase and DFS features interact and merge, we use group normalization to achieve the attention of different channels. The generated attention maps are thresholded to get the strengthened and weakened maps, and then the strengthened and weakened Phase and DFS features are obtained. Based on retaining important features and attenuating minor features  [43], the feature space redundancy is reduced while utilizing both features.
@@
Separate operation aims to separate those informative feature maps from less informative ones corresponding to the spatial content. We leverage the scaling factors in Group Normalization (GN)  [44] layers to assess the informative content of different feature maps. To be concrete, given an intermediate feature map $X_{\text{PD}}\in\mathbb{R}^{N\times C\times H\times W}$, where $N$ is the batch axis, $C$ is the channel axis, $H$ and $W$ are the spatial height and width axes. We first standardize the input feature $X$ by subtracting mean $\mu$ and dividing by standard deviation $\sigma$ as follows:
@@
$X_{\text{gn}}=\text{GN}(\mathbf{X}_{\text{PD}})=\gamma\frac{X_{\text{PD}}-\mu}{\sqrt{\sigma^{2}+\epsilon}}+\beta,$ (13)
@@
where $\mu$ and $\sigma$ are the mean and standard deviation of $X_{\text{PD}}$, $\epsilon$ is a small positive constant added for numerical stability, $\gamma$ and $\beta$ are trainable affine transformation parameters.
@@
Noted that we leverage the trainable parameters $\gamma\in\mathbb{R}^{C}$ in GN layers as a way to measure the variance of spatial pixels for phase and DFS. The richer spatial information reflects more variation in spatial pixels contributing to a larger $\gamma$. The normalized correlation weights $W_{\gamma}\in\mathbb{R}^{C}$ are obtained by equation 14, which indicates the importance of different phase and DFS feature maps.
$W_{\gamma}=\{w_{i}\}=\frac{\gamma_{i}}{\sum_{j=1}^{C}\gamma_{j}},\quad i,j=1,2,\dots,C.$ (14)
@@
Then the weight values of feature maps reweighted by $W_{\gamma}$ are mapped to the range (0, 1) by the sigmoid function and gated by a threshold. We set those weights above the threshold to 1 to obtain the informative weights $W_{1}$ while setting them to 0 to gain the non-informative weights $W_{2}$ (the threshold is set to 0.5 in the experiments). The whole process of acquiring $W$ can be expressed as equation (16):
$W=Gate(\text{Sigmoid}(W_{\gamma}(GN(X_{\text{PD}})))).$ (15)

### v4

# Wi-CBR: Salient-aware Adaptive WiFi Sensing for Cross-domain Behavior Recognition

@@ Abstract
The challenge in WiFi-based cross-domain Behavior Recognition lies in the significant interference of domain-specific signals on gesture variation. However, previous methods alleviate this interference by mapping the phase from multiple domains into a common feature space. If the Doppler Frequency Shift (DFS) signal is used to dynamically supplement the phase features to achieve better generalization, it enables the model to not only explore a wider feature space but also to avoid potential degradation of gesture semantic information. Specifically, we propose a novel Salient-aware Adaptive WiFi Sensing for Cross-domain Behavior Recognition (Wi-CBR), which constructs a dual-branch self-attention module that captures temporal features from phase information reflecting dynamic path length variations while extracting kinematic features from DFS correlated with motion velocity. Moreover, we design a Saliency Guidance Module that employs group attention mechanisms to mine critical activity features and utilizes gating mechanisms to optimize information entropy, facilitating feature fusion and enabling effective interaction between salient and non-salient behavioral characteristics. Extensive experiments on two large-scale public datasets (Widar3.0 and XRF55) demonstrate the superior performance of our method in both in-domain and cross-domain scenarios.
Codehttps://github.com/zrbwsw/Wi-CBR
Introduction
Human Behavior Recognition (HBR) enables systems to intelligently interpret human actions and is widely applied in areas like human-computer interaction (tang2025discrete; tang2025sign), Interactive integration (song2024emotional; song2023emotion), action understanding (xu2025towards; zhang2025temporal), and accessible communication (tangslt2022; tang2022gloss). WiFi-based HBR, leveraging Channel State Information (CSI) and Received Signal Strength Indicator (RSSI), has gained attention for its ability to model motion-induced signal changes (1data2; 1data3). While coarse-grained activities are easily detectable, fine-grained gestures remain challenging due to subtle movements (14). By mapping CSI variations to behaviors, WiFi sensing supports applications in homes, workplaces, and other indoor environments (Wigest; WiFinger).
Figure 1: When the user and behavior remain unchanged, relative to the location of transmitter–receiver (Tx-Rx) pairs or the overall environment changes, the phase changes significantly, and DFS has a certain degree of domain independence.
An important challenge in WiFi-based behavior recognition is environmental dependence, as wireless signals often carry environment-specific information that is unrelated to human behavior(zhang2025wiopen). Cross-domain recognition accuracy significantly decreases due to differences in data distribution, noise, and changes in the propagation path. Traditional methods are divided into Modeling-Based and Learning-Based approaches (2modelandlearn1; 2modelandlearn2). Modeling-Based methods use handcrafted features that are robust to environmental variation but lose raw signal richness, limiting performance in complex tasks. Learning-Based methods leverage deep neural networks to learn directly from raw CSI but are hindered by environmental sensitivity and a lack of domain-invariant guidance. Therefore, how to integrate original features with domain-independent features is key to achieving effective cross-domain generalization (guo2024style; zhang2025style; wu2025percept; yin2022mix; guo2025domain), bridging the gap between controlled training environments and real-world applications.
Our observations find that, in wireless communication systems, the characteristics of the received signal are influenced by various factors, including the positions of the transmitter (Tx) and receiver (Rx), the transmission path, and the dynamics of the overall environment. As shown in Fig. 1, significant variations in phase values occur when the receiver’s position changes, even if the user and behavior remain constant. Figure (b) further demonstrates that phase values are highly sensitive to environmental changes, exhibiting unstable patterns across different environments and receiver positions. In contrast, the DFS shows greater domain independence and stability, with its features generally remaining consistent across varying environments and receiver locations. Building on these observations, the proposed method leverages the stability of DFS features while addressing the sensitivity of phase values. By extracting and aligning domain-invariant DFS features, the approach enhances the adaptability to dynamic and multi-domain scenarios. Furthermore, we aim to preserve the rich temporal-spatial patterns of raw signals while embedding domain-independent priors that facilitate robust cross-domain generalization. This hybrid motivation directly informs the network design, enabling complementary feature learning and improving cross-domain performance.
@@ Introduction
In this work, we propose Wi-CBR to leverage the full spectrum of raw data while ensuring the integration of domain-independent features. Specifically, our method introduces a novel multimodal collaborative awareness framework that efficiently combines phase data, which captures dynamic path length changes, and DFS data, which reflects frequency shifts tied to gesture speed. To capture spatial-temporal patterns within each modality, we employ a Two-branch self-attention module, enabling the system to focus on important temporal and spatial features within each signal type. A group attention mechanism is then applied to the concatenated phase and DFS features, allowing the model to identify key group features that are essential for behavior recognition. Finally, a gating mechanism is used to divide the fused features into enhancement and suppression branches, optimizing information entropy and facilitating collaborative complementarity. This fusion of multiple data sources, along with the innovative use of attention and gating mechanisms, enables more accurate and robust behavior recognition, particularly in cross-domain scenarios. Our main contributions are summarized as follows:
  • We propose Wi-CBR, a novel Salient-aware Adaptive WiFi Sensing framework for Cross-domain Behavior Recognition. Wi-CBR proposes a two-branch self-attention module that captures temporal features from phase information, reflecting dynamic path length variations, while extracting kinematic features from DFS, which are correlated with motion velocity.
  • We design a Saliency Guidance Module that leverages group attention mechanisms to identify critical activity features. This module incorporates gating mechanisms to optimize information entropy, facilitating feature fusion and enabling effective interaction between salient and non-salient behavioral characteristics.
  • Extensive experiments on two large-scale public datasets demonstrate the superior performance of our method in both in-domain and cross-domain scenarios.
Related Work
Modeling-Based HBR
Modeling-based approaches preprocess raw CSI data, extract manual features (e.g., velocity statistics), and use machine learning for gesture recognition (WiFinger; WIHGR; 2model6; 2model3; 2model2; Wi-NN). WiGest (Wigest) relies on coarse-grained RSS, limiting accuracy. WiMU (WiMU) struggles with scalability, and WiDraw (Widraw) requires over 25 transceivers, making it impractical. QGesture (Qgesture) uses two antennas but depends on prior hand position knowledge. (2model3) introduces dynamic phase exponential error for gesture quality, while Wi-NN(Wi-NN) applies time-domain feature selection with KNN classification. These methods link WiFi signals to gestures but neglect environmental impact, as gesture performance in varying environments alters WiFi waveforms (Chen2023; Kang2021). Widar 3.0 (Widar3.0) introduces domain-agnostic BVP features, and WiHF (WIHF) derives domain-independent motion patterns; however, handcrafted features limit the capture of spatiotemporal cues. WiGesture (WiGesture) focuses on position-independent Motion Navigation Primitives (MNP). WiGNN(WiGNN) focuses on graph modeling for multi-receiver topologies through GNN-based temporal-frequency aggregation and data augmentation.
Learning-Based HBR
Learning-based methods directly process raw CSI data, such as amplitude and phase, for automatic pattern recognition. Wikey (Wikey) enables keystroke recognition but is highly sensitive to environmental changes. WiSign (WiSign) extracts spatio-temporal features for sign language recognition but requires extensive domain-specific training. Tong et al.(Tong2023) introduced a CNN-GRU-Attention (CGA) model with phase correction and gesture truncation for improved data validity, while Yang et al.(Yang2019) proposed a CNN-RNN architecture for enhanced spatiotemporal pattern learning. WiHGR (WIHGR) uses a phase difference matrix and an improved ABGRU for feature extraction. To address cross-domain challenges, CROSSGR (CrossGR) extracts gesture-related features independent of users. WiGr (WiGr) uses query-class prototype similarity to mitigate cross-domain variations. WIGRUNT (WIGRUNT) applies a spatio-temporal dual-attention network with ResNet for feature extraction, while Wi-SFDAGR (Wi-SFDAGR) addresses cross-domain issues using Unsupervised Domain Adaptation (UDA) for unlabeled test data scenarios
Method
Figure 2: Framework Overview of Wi-CBR. The proposed Wi-CBR framework integrates Signal Preprocessing, a Two-branch Self-attention Module, and a Saliency Guidance Module to achieve robust behavior recognition across domains. It leverages phase information for temporal dynamics and Doppler Frequency Shift (DFS) for spatial motion characteristics. A Group Attention Interaction mechanism identifies salient features, while Saliency Adaptive Fusion effectively combines critical and non-critical behavior features. The system is designed to generalize across different environments, orientations, and locations, ensuring superior performance in both in-domain and cross-domain scenarios.
@@
As shown in Fig. 2, Wi-CBR consists of four components: signal preprocessing, two-branch self-attention learning, saliency guidance, and classification prediction. In signal preprocessing, the CSI-ratio model handles denoising, while DFS is extracted using STFT. The phase and DFS matrices are visualized as images for deep learning. The network employs two-branch self-attention and pre-trained ResNet18 for initial feature extraction, followed by feature fusion via a cross-model interactive module. Finally, behavior prediction is achieved using a classifier with dual-loss constraints.
Task Definition
@@
$H(f,t)$ $=\sum_{m_{s}\in\Phi_{s}}a_{m_{s}}(f,t)e^{-j2\pi\frac{d_{m_{s}}(t)}{\lambda}}$
$+\sum_{m_{d}\in\Phi_{d}}a_{m_{d}}(f,t)e^{-j2\pi\frac{d_{m_{d}}(t)}{\lambda}}$ (3)
CSI Denoising Preprocessing
@@
As demonstrated in the previous section, the gesture can be portrayed by the change of phase shift in CSI. Unfortunately, for commodity WiFi devices, as the transmitter and receiver are not synchronized, there exists a time-varying random phase offset $e^{-j\theta_{\text{offset}}}$ :
$H(f,t)$ $=e^{-j\theta_{\text{offset}}}(a_{m_{s}}(f,t)e^{-j2\pi\frac{d_{m_{s}}(t)}{\lambda}}$
$+a_{m_{d}}(f,t)e^{-j2\pi\frac{d_{m_{d}}(t)}{\lambda}}),$ (4)
@@
where $e^{-j2\pi\frac{d(t)}{\lambda}}$, and $d(t)$ denote phase-shift, and path length of dynamic components, respectively. This random offset thus prevents us from directly using the CSI phase information.
@@
Therefore, we need to eliminate $e^{-j\theta_{\text{offset}}}$. Fortunately, for commodity WiFi cards, this random offset remains the same across different antennas on the same WiFi network interface card (NIC) as they share the same RF oscillator. Thus, it can be eliminated by the CSI-ratio model:
$H_{r}(f,t)=\frac{H_{1}(f,t)}{H_{2}(f,t)}=\frac{e^{-j\theta_{\text{offset}}}(H_{s,1}+H_{d,1})}{e^{-j\theta_{\text{offset}}}(H_{s,2}+H_{d,2})}$ (5)
@@
where $H_{1}(f,t)$ and $H_{2}(f,t)$ are the CSI of two receiving antennas. When two antennas are close to each other, $\Delta d$ can be regarded as a constant. According to Mobius transformation  (csi1), (5) represents transformations such as scaling and rotation of the phase-shift $e^{-j2\pi\frac{d_{1}(t)}{\lambda}}$ of antenna 1 in the complex plane, and these transformations will not affect the changing trend of the phase-shift(csi2; csi3; csi4).
CSI to Phase. Phase ratio P extracted from $H_{r}$ can be used to describe gestures:
$\text{P}=angle(H_{r}),$ (6)
@@
where $angle(\cdot)$ denotes the phase extraction function. For a complex $z=\text{abs}(z)\cdot e^{j\theta}$, we can use $angle(\cdot)$ to obtain the phase of $z$, $\theta=angle(z)$.
CSI to DFS. To extract Doppler Frequency Shift (DFS) features, we take the raw CSI signal received by a specific Tx–Rx antenna pair, denoted as $H_{q}(f,t)$, as an example. This stream contains temporal variations caused by human motion. Unlike $H_{r}(f,t)$, which is a ratio of two antenna CSI streams designed to remove phase offset for phase analysis, $H_{q}(f,t)$ retains the original frequency content, making it more suitable for Doppler-based gesture analysis. The DFS reflects frequency changes due to gesture motion, such as hand speed or direction.
$S_{q}(\tau,\omega)=\int_{-\infty}^{\infty}H_{q}(f,t)w(t-\tau)e^{-j\omega t}dt\ ,$ (7)
@@
where $w(t-\tau)$ is a window function (e.g., Hanning window) to segment the signal, $\tau$ is the time localization, and $\omega$ is the frequency in radians, tied to the Doppler shift. The power spectrogram is then computed:
$\text{D}=|S_{q}(\tau,\omega)|^{2}=\left|\int_{-\infty}^{\infty}H_{q}(f,t)w(t-\tau)e^{-j\omega t}dt\right|^{2}\$(8)
@@
In this spectrogram, the frequency $\omega$ at each time $\tau$ corresponds to the Doppler shift $f_{d}=\omega/(2\pi)$. $f_{d}$ is proportional to the gesture speed $v(t)$. Positive frequencies indicate motion toward the receiver, while negative frequencies indicate motion away. This time-frequency representation enables detailed analysis of gesture dynamics.
For the Widar3.0 dataset, each DFS file is a $6\times 121\times\mathrm{T}$ matrix, where the first dimension represents the 6 receivers, the second dimension represents the 121 frequency segments ranging from [60,60][-60,60] Hz, and the third dimension represents the timestamps with a sampling rate of 1000 Hz.
Two-branch Self-attention Learning
@@
$\mathbf{A}_{p}$ $=\sigma\left(\text{BN}\left(f_{\text{conv}}^{(7)}\left(f_{\text{conv}}^{(7)}(\mathbf{P}_{x})\right)\right)\right).$ (9)
where $\sigma(\cdot)$ is the sigmoid function. $\mathbf{A}_{d}$ is computed in the same way using $\mathbf{D}_{x}$ as input. The refined features are obtained through:
$\mathbf{P}_{\text{out}}$ $=\mathbf{P}_{x}\otimes\mathbf{A}_{p}\oplus\mathbf{P}_{x};$ $\quad\mathbf{D}_{\text{out}}$ $=\mathbf{D}_{x}\otimes\mathbf{A}_{d}\oplus\mathbf{D}_{x}$ (10)
where $\otimes$ is element-wise multiplication, $\oplus$ is element-wise summation. Each branch maintains independent convolution parameters, with channel dimensions preserved in the first convolution ($3\rightarrow 3$) and reduced to a single channel ($3\rightarrow 1$) in the second convolution.
@@
$\mathbf{X}_{\text{PD}}=\mathrm{Concat}(\phi_{\text{ResNet}}^{P}(\mathbf{P}_{\text{out}}),\phi_{\text{ResNet}}^{D}(\mathbf{D}_{\text{out}}))\in\mathbb{R}^{1024\times 7\times 7}$ (11)
Saliency Guidance Module
@@
Group Attention Interaction. To make the extracted Phase and DFS features interact and merge, we use group normalization to achieve the attention of different channels. The generated attention maps are thresholded to get the strengthened and weakened maps, and then the strengthened and weakened Phase and DFS features are obtained. Based on retaining important features and attenuating minor features  (scconv), the feature space redundancy is reduced while utilizing both features.
@@
Separate operation aims to separate those informative feature maps from less informative ones corresponding to the spatial content. We leverage the scaling factors in Group Normalization (GN)  (GN) layers to assess the informative content of different feature maps. To be concrete, given an intermediate feature map $X_{\text{PD}}\in\mathbb{R}^{N\times C\times H\times W}$, where $N$ is the batch axis, $C$ is the channel axis, $H$ and $W$ are the spatial height and width axes. We first standardize the input feature $X$ by subtracting mean $\mu$ and dividing by standard deviation $\sigma$ as follows:
@@
$X_{\text{gn}}=\mathrm{GN}(\mathbf{X}_{\text{PD}})=\gamma\frac{X_{\text{PD}}-\mu}{\sqrt{\sigma^{2}+\epsilon}}+z$ (12)
@@
where $\mu$ and $\sigma$ are the mean and standard deviation of $X_{\text{PD}}$, $\epsilon$ is a small positive constant added for numerical stability, $\gamma$ and $z$ are trainable affine transformation parameters.
@@
Noted that we leverage the trainable parameters $\gamma\in\mathbb{R}^{C}$ in GN layers as a way to measure the variance of spatial pixels for phase and DFS. The richer spatial information reflects more variation in spatial pixels contributing to a larger $\gamma$. The normalized correlation weights $W_{\gamma}\in\mathbb{R}^{C}$ are obtained by equation (14), which indicates the importance of different phase and DFS feature maps.
$W_{\gamma}=\{w_{i}\}=\frac{\gamma_{i}}{\sum_{j=1}^{C}\gamma_{j}},\quad i,j=1,2,\dots,C$ (13)
@@
Then the weight values of feature maps reweighted by $W_{\gamma}$ are mapped to the range (0, 1) by the sigmoid function and gated by a threshold. We set those weights above the threshold to 1 to obtain the informative weights $G_{1}$ while setting them to 0 to gain the non-informative weights $G_{2}$ (the threshold is set to 0.5 in the experiments):
$G=\text{Gate}(\text{Sigmoid}(W_{\gamma}(GN(X_{\text{PD}}))))$ (14)